OpenStudy (anonymous):

5 , 2 , -1 , -4, -7 , . . . express the terms using explicit formula

OpenStudy (anonymous):

an = a1 + (n-1)d d=-3 and a1 = 5 an = 5 + (n-1)(-3) = 8 - 3n